When a 4.00 g sample of KCN is dissolved in water in a calorimeter that has a total heat capacity of 1.94 kJ⋅K−1, the temperature decreases by 0.370 K. Calculate the molar heat of solution of KCN.
1
Expert's answer
2019-10-22T05:44:13-0400

Moles of "KCN=\frac{4}{65}=0.0615mol"

Total heat capacity"=n\times C\times∆T"

"1.94=0.0615\times C\times0.370"

"C=85.25\frac{KJ}{mol}"